A rectangular solid has a base with length 4cm and width 7cm. If the volume of the solid is 224cm^3, find the height of the solid. (Hint: The volume of a rectangular solid is given by V=LWH)

Answer by checkley71(8403)

VOLUME=L*W*H
224=4*7*H
224=28H
H=224/28
H=8 CM IS THE HEIGHT.
PROOF
224=4*7*8
224=224